OpenStudy (princesssleelee):

Which graph represents the function of f(x) = the quantity of 9 x squared plus 9 x minus 18, all over 3 x plus 6? graph of 3 x plus 6, with discontinuity at 2, 12 graph of 3 x plus 6, with discontinuity at negative 2, 0 graph of 3 x minus 3, with discontinuity at negative 2, negative 9 graph of 3 x minus 3, with discontinuity at 2, 3

OpenStudy (aihberkhan):

One moment, let me see.. :)

OpenStudy (tkhunny):

Show YOUR work, first. Let's see what YOU get - sure or not.

OpenStudy (aihberkhan):

Yes please do.

OpenStudy (princesssleelee):

i have no work on this one, i dont know how to even start! lol

OpenStudy (aihberkhan):

Okay well is this what you mean? \[f(x) = \frac{ 9x^2 + 9x - 18 }{ 3x + 6 }\]

OpenStudy (tkhunny):

Insufficient response. Some notational improvement would also be helpful. f(x) = the quantity of 9 x squared plus 9 x minus 18, all over 3 x plus 6? This is a good representation. f(x) = (9x^2 + 9x - 18)/(3x+6) With a little LaTex, you can write it as Aihber has suggested. Please demonstrate that you can factor the Denominator.

OpenStudy (aihberkhan):

Well to start you off, so you can at least START... try factoring on your own. :) Can you do that? @princesssleelee

OpenStudy (princesssleelee):

i do not know how

OpenStudy (aihberkhan):

Okay well lets start with \(3x\) and \(6\)

OpenStudy (aihberkhan):

What is a common factor between those two?

OpenStudy (princesssleelee):

3

OpenStudy (aihberkhan):

Right! So can you try and rewrite \(3x + 6\) while pulling out a 3?

OpenStudy (aihberkhan):

Hmm... I know the way I said it was a bit confusing. So let me explain further.

OpenStudy (tkhunny):

You MUST be able to factor the Denominator. You CANNOT solve this problem without these basic skills. If you have been given this problem without ANY exposure to such factoring, then whoever gave you the problem is an incompetent clod. Hoping to believe that your teacher is more capable than that, I choose to believe that you CAN factor the denominator.

OpenStudy (princesssleelee):

3x-3 is the entire function factored, correct?

OpenStudy (aihberkhan):

I appreciate you trying! :) Unfortunately, that is incorrect. Also, it is better if you are able to use parenthesis. Also, why did you get a minus sign instead of a plus sign? Can you try again?

OpenStudy (princesssleelee):

then i have no idea how to do it

OpenStudy (tkhunny):

That's the spirit. Look really hard at the function of the Distributive Property: a(b+c) = ab + ac It goes the other way, too. de + df = d(e+f) Do that with (3x+6). Find a factor that is common when comparing '3x' and '6'.

OpenStudy (princesssleelee):

this is just so confusing

OpenStudy (aihberkhan):

Yes, I know. It is a bit confusing. Well, I am not sure why you changed \(3x + 6\) into a subtraction problem. It should stay positive. So other than that you are correct. Instead of \(3x-2\) it would be \(3(x + 2)\)

OpenStudy (aihberkhan):

Do you see that? :)

OpenStudy (tkhunny):

Note: If you cannot factor the denominator, you have no prayer of factoring the numerator, which is your next task. Just look at the rule and apply it. It doesn't have to be confusing. Waiting for it to bite you is a little confusing. Logically address the rule. Logically address the given problem statement.

OpenStudy (aihberkhan):

Right :)

OpenStudy (aihberkhan):

But, do you see how I did that? @princesssleelee

OpenStudy (aihberkhan):

Take your time to evaluate.

OpenStudy (princesssleelee):

yes i do for the most part.

OpenStudy (princesssleelee):

then distribute?

OpenStudy (aihberkhan):

No you would not distribute. If you did, you would get the same thing as before. Factoring is basically undoing the fact that you distributed. So, now we do the same with the numerator. Do you think you can factor \(9x^2 + 9x -18\)?

OpenStudy (princesssleelee):

well they all go into 9

OpenStudy (princesssleelee):

x2+x-2

OpenStudy (aihberkhan):

Right!! \[x^2 + x - 2\]

OpenStudy (aihberkhan):

Now, what are two integer numbers that multiply to be \(-2\) and add up to be \(1\)?

OpenStudy (aihberkhan):

Because, we must continue factoring.

OpenStudy (princesssleelee):

1 and -2

OpenStudy (princesssleelee):

nevermind thats wrong

OpenStudy (aihberkhan):

You have the numbers right but the positive and negative signs mixed up. Can you try again?

OpenStudy (aihberkhan):

Yes, try again :)

OpenStudy (princesssleelee):

2 and -1

OpenStudy (aihberkhan):

Right! So the factors would be: \((x + 2)\) and \((x - 1)\)

OpenStudy (aihberkhan):

Do you see that?

OpenStudy (princesssleelee):

yes

OpenStudy (aihberkhan):

Okay so, \(9x^2+9x-18\) is now \(9(x+2)(x-1)\)

OpenStudy (aihberkhan):

So it would completely be: \[f(x) = \frac{ 9(x + 2)(x - 1) }{ 3(x + 2) }\]

OpenStudy (aihberkhan):

Do you see that?

OpenStudy (princesssleelee):

yes

OpenStudy (aihberkhan):

We can't divide by 0 meaning that \(x+2\) can't be 0. However, \(x+2\) can be 0 when \(x\) is... what? Do you know? @princesssleelee

OpenStudy (princesssleelee):

-2

OpenStudy (aihberkhan):

Right so \[f(x) = \frac{ 9(x + 2)(x - 1) }{ 3(x + 2) } = 3(x - 1), x \neq -2\]

OpenStudy (aihberkhan):

Do you see that?

OpenStudy (aihberkhan):

This just means that our function will not exist at \(x = -2\)

OpenStudy (princesssleelee):

ok

OpenStudy (aihberkhan):

So... if we have \(f(x)=3x-3\), then the y-intercept is -3. Right? We also know that \(f(-2)=3(-2)-3=-6-3=-9\) Now when graphing... we would draw a line going through \((0,-3)\) and \((-2,-9)\) However, we will erase the one point at (-2,-9) since our function shouldn't exist there

OpenStudy (aihberkhan):

So the graph should look something like this: